OpenStudy (anonymous):

find the exact amplitude, period, and horizontal shift of 2tan(pix+3) I have amplitude = 2 period = 2

jimthompson5910 (jim_thompson5910):

The period of y = A*tan(Bx+C) + D is T = pi/B In your case, B = pi, so the period is T = pi/pi = 1

jimthompson5910 (jim_thompson5910):

For y = A*tan(Bx+C) + D, the value of C determines the horizontal shift if C is positive, then the graph of y = A*tan(Bx) + D is shifted C units to the left if C is negative, then the graph of y = A*tan(Bx) + D is shifted C units to the right

OpenStudy (anonymous):

so its shifted 3 units to the left?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np